# font=itshape, raggedright, noindent, leftmargin=0.08in, rightmargin=0.08in, begintext=“ , endtext= ”

### v2

@@ 1 Introduction
To respond to this inquiry, we study the implicit data distribution favored by trained models. In other words, our approach to understanding a model is based on creating samples in the data domain that the trained model considers favorable for a specific task. Unlike conventional ML pipelines that focus on static datasets and predictive accuracy, our approach enables dynamic interrogation of model behavior via investigating the answers of the model to specific questions. We show that these questions can be customized to each situation and they can be expressed mathematically through a loss function that evaluates the data based on a combination of data characteristics and model parameters. We consider the problem of understanding a model to be a more nuanced endeavor that requires exploration across multiple dimensions of questioning. This involves providing explanations, such as counterfactual [28] or prototypical [3] scenarios, shedding light not only on why a particular prediction was made, but going beyond it as well. For instance, insights into model behavior can be gained by generating parameter-sensitive data samples. When two models showing similar performances give different predictions, also known as predictive multiplicity [22], our approach can be used to generate data to systematically compare their behaviors in diverse scenarios. These custom questions, and others, provide a qualitative understanding of the model. In addition, users have the flexibility to customize queries by designing specific probing functions within the data space.
@@
Related Literature. Our work complements extensive research in synthetic data generation that has been pivotal in addressing fairness, bias detection [15] and reduction [30, 27] as well as dataset augmentation [29, 6]. Generative Adversarial Networks (GANs) and Variational Autoencoders (VAEs) have been widely used to approximate original data distributions [11, 30, 14, 4], focusing on privacy, diversity, and fidelity as primary goals.
@@
Recent studies leveraged generative models for counterfactual generation and exploring underrepresented data regions. For example, Joshi et al. [12] proposed a framework for generating task-specific synthetic data, enhancing model explainability. Similarly, Redelmeier et al. [25] introduced an approach using autoregressive generative models to create counterfactuals, facilitating bias exploration and decision boundary analysis. Recent work on global counterfactual explanations has further expanded the scope of interpretability by targeting group-level understanding. Rawal and Lakkaraju [24] introduced a framework for generating global, rule-based recourse summaries for population subgroups, optimizing objectives like accuracy, coverage, and cost. These summaries are ‘if-then’ rules linking subgroup characteristics to actionable feature changes that influence a model’s prediction. Plumb et al. [23] proposed interpreting clusters in low-dimensional representations by finding sparse transformations that align one group with another.
@@
Energy-based models (EBMs) have also emerged as a promising framework, combining generative and discriminative modeling tasks. By treating classifier logits as an energy function, EBMs can model joint distributions over data and labels [17, 5]. Applications of EBMs include adversarial robustness, out-of-distribution detection, and data augmentation [31, 19, 1, 21]. For instance, Duvenaud et al. [5] demonstrated improved out-of-distribution detection using a joint energy-based model, while Ma et al. [20] extended EBMs to tabular data for synthetic data generation.
@@
The proposed framework draws inspiration from these works while introducing a distinct perspective. Our probing function can be seen as an energy function and leads to Gibbs distribution. However, rather than learning the energy function to capture the data distribution (conditioned on label), we create a probing function using trained models. This design allows the distribution to generate samples that address the specific posed question. Related works, such as [5] and [20], adopt a similar approach by utilizing a trained classifier to obtain an energy function and using Langevin dynamics for sampling from the Gibbs distribution. However, their main objective is to mimic the true data distribution. In fact, the former paper combines training of the energy function and classifier. In contrast, we propose a flexible framework that allows for directing diverse queries to trained models via probing functions that reflect various objectives, such as identifying prediction-risky, parameter-sensitive, or model-contrastive data samples.
Contributions. We contribute to the literature by introducing a novel inductive approach. This approach creates data samples using a probing function that engages a trained model. Through this framework, the generated samples provide answers to various questions posed to trained models. With a computational study, we support our approach by applying it to a range of classification and regression tasks, showcasing its effectiveness in generating data tailored to specific queries. Our numerical results demonstrate the flexibility of our approach in uncovering biases, facilitating model interpretability, and consequently, promoting alignment of model predictions and human preferences.
@@ 2 The Mathematical Framework
First, our notation: The labeled data lie in $\mathcal{X}\times\mathcal{Y}$, and the model defines a predictor function $f({\boldsymbol{\mathbf{\theta}}},\cdot):\mathcal{X}\to\mathcal{Y}^{\prime}$ for any given set of model parameters ${\boldsymbol{\mathbf{\theta}}}\in\Theta$. For a given sample ${\boldsymbol{\mathbf{x}}}\in\mathcal{X}$, the predicted label $y_{{\boldsymbol{\mathbf{\theta}}}}({\boldsymbol{\mathbf{x}}})\in\mathcal{Y}$ is obtained from the predictor function. The cost function $\ell_{F}:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$ measures how far the predicted labels are from the true labels.
@@
Figure 1: (left) Overview of model probing by data generation. Samples drawn from $p^{*}({\boldsymbol{\mathbf{x}}})$ constitute an answer to the question posed by $G$. The vertical arrows (a) and (c) start with functions and lead to distributions on the same space by solving the Bayesian Learning Problems (2) and (3). The diagonal arrow (b) obtains a loss function on the data space by integrating out the ${\boldsymbol{\mathbf{\theta}}}$ dependence from a designed function via $q^{*}({\boldsymbol{\mathbf{\theta}}})$. (right) Specializing to a linear regression model with mean square error and no regularizer admits an analytic solution. The distribution of data points ${\boldsymbol{\mathbf{x}}}$ which are preferred by the LR model in giving a chosen prediction $y^{\prime}$ is calculated to be a Gaussian distribution centered at a point $\hat{{\boldsymbol{\mathbf{f}}}}$, which is shifted from the mean of given data by a certain amount depending on the desired output value $y^{\prime}$. Explicit forms of $\hat{{\boldsymbol{\mathbf{f}}}},\Sigma$ and the derivation can be seen in Appendix A.
The standard construction of the parameter loss function is
$F({\boldsymbol{\mathbf{\theta}}})=\int_{\mathcal{X}\times\mathcal{Y}}(\ell_{F}(y_{\boldsymbol{\theta}}(\mathbf{x}),y)+R_{F}(\boldsymbol{\theta}))\mathrm{d}\nu(\mathbf{x},y)=\frac{1}{N}\sum_{i=1}^{N}\ell_{F}(y_{{\boldsymbol{\mathbf{\theta}}}}({\boldsymbol{\mathbf{x}}}_{i}),y_{i})+R_{F}({\boldsymbol{\mathbf{\theta}}}),$ (1)
which can be seen as an integral of $\ell_{F}+R_{F}$ against the empirical distribution given by the training dataset $\{({\boldsymbol{\mathbf{x}}}_{i},y_{i})\}_{i=1}^{N}\subseteq\mathcal{X}\times\mathcal{Y}$. Here, $R_{F}({\boldsymbol{\mathbf{\theta}}})$ is a regularizer term that depends only on ${\boldsymbol{\mathbf{\theta}}}$.
We propose a framework for probing a model with answers in the form of generated data and it is structured to be symmetric to the training process itself: instead of generating parameters given a data distribution, we generate data given a trained model parameter distribution. Similar to how we form the loss $F$ on the parameters, our data loss $G$ can be obtained by integrating out the ${\boldsymbol{\mathbf{\theta}}}$ dependence of a function that depends on both the data and the model parameters. We design this latter function to be low when the desiderata required by a probing scenario are satisfied. Depending on our choice, we end up probing the model with different questions and gain valuable insights into the trained model’s behavior by observing the generated data, both qualitatively and via its population statistics. Figure 1 presents the overview of our framework.
The blue arrows (a) and (c) in Figure 1 represent solving the Bayesian Learning Problem (BLP), which –instead of finding a single set of model parameters ${\boldsymbol{\mathbf{\theta}}}^{*}$– seeks a distribution that balances the tasks of minimizing the expected loss and maximizing the entropy. In case of (a), we solve
@@
where $\mathcal{Q}$ is a choice of candidate distributions on $\Theta$, and $\mathcal{H}(q)=-\int_{\Theta}q\log q\,\mathrm{d}\mu$ is the entropy with respect to a base measure $\mu$. The problem can be interpreted as an implementation of the exploration-exploitation trade-off in the parameter space. The constant $\tau>0$, called the temperature, controls the balance between these two objectives. Absent any restriction, i.e., if $\mathcal{Q}$ is the set of all probability distributions [^1] All probability distributions which are absolutely continuous with respect to a given base measure, which in this case taken to be the Lebesgue measure $\mathrm{d}\,{\boldsymbol{\mathbf{\theta}}}$., then the Gibbs-Boltzmann distribution $q^{*}({\boldsymbol{\mathbf{\theta}}})\propto e^{-\frac{1}{\tau}F({\boldsymbol{\mathbf{\theta}}})}$ is the unique solution to (2).
Standard stochastic gradient descent can be seen as a specialization of this setup. Given a very restrictive family $\mathcal{Q}$, such as the manifold of Dirac delta distributions [^2] To eschew technicalities of infinities, instead of exactly using Dirac delta we may instead consider distributions which are supported everywhere, and highly concentrated around a point but with fixed variance. supported on a single ${\boldsymbol{\mathbf{\theta}}}$, the entropy term becomes irrelevant, and we get the classical optimization problem of minimizing $F$.
Symmetrically on the data space $\mathcal{X}$, the blue arrow labeled (c) in Figure 1, represents solving
@@
The expectation term encourages the solution $p^{*}({\boldsymbol{\mathbf{x}}})$ to concentrate its mass on data samples where $G$ is minimized, but the entropy term encourages $p^{*}({\boldsymbol{\mathbf{x}}})$ to explore a wide variety of data samples. Just like $F$ in (1), the function $G$ is constructed as an integral of a function of two ${\boldsymbol{\mathbf{\theta}}}$ and data against the distribution $q^{*}({\boldsymbol{\mathbf{\theta}}})$. Training in the parameter space encourages finding those parameters where $F$ is lower, so the reached parameters are more compatible with the given training data. Similarly, training with the loss function $G$ in the data space with (3) encourages finding data points which are compatible with the trained parameters sampled from $q^{*}$ (which may be a single set of parameters $\theta^{*}$ in case it is a Dirac-delta distribution). In this way, the samples from the distribution $p^{*}({\boldsymbol{\mathbf{x}}})$ are considered to be an answer to the question put forth by $G$: Which data points makes the $G$-loss low while maintaining diversity among the produced samples? Then, it remains to design $G$ so that this coincides with illuminating questions about the trained model.
We may explicitly select $\mathcal{P}$ as a family of distributions depending on the nature of the data distribution. Alternatively, we can retain the full space of probability distributions but replace $G$ with $G\circ\varphi$ for some function $\varphi:\mathcal{Z}\to\mathcal{X}\times\mathcal{Y}$. In the latter case, $\mathcal{Z}$ can be the latent space and $\varphi$ can be the decoder function of a trained variational autoencoder. This procedure produces a distribution on $\mathcal{Z}$, which we can sample from and map to $\mathcal{X}$ via $\varphi$, thus sampling data from the pushforward distribution.
There are various methods of solving (3). In the examples given below, we used Metropolis Adjusted Langevin Algorithm (MALA) to sample directly from the Gibbs-Boltzmann distribution $p^{*}({\boldsymbol{\mathbf{x}}})\propto e^{{-\frac{1}{\tau}G({\boldsymbol{\mathbf{x}}})}}$. Details of this method are given in Appendix C. Instead of MALA, one can also apply Picard Iterations to solve the overdamped Langevin stochastic differential equation, or alternatively apply Variational Inference (VI) to reach a distribution $p^{*}$; see [8, 9].
@@ 3 Probing Trained Models
We start with a general structure of the loss function for questioning trained models:
$G({\boldsymbol{\mathbf{x}}})=\int_{\Theta}\ell_{G}(y_{{\boldsymbol{\mathbf{\theta}}}}({\boldsymbol{\mathbf{x}}}),\widehat{y}({\boldsymbol{\mathbf{x}}}))q^{*}({\boldsymbol{\mathbf{\theta}}})\mathrm{d}\,{\boldsymbol{\mathbf{\theta}}}+R_{G}({\boldsymbol{\mathbf{x}}}),$ (4)
@@
Figure 2: Given a dataset of two concentric circles labeled red and blue, two Support Vector Machine (SVM) models are trained on the binary classification task with kernels chosen as Radial Basis Function (RBF) and cubic polynomial, respectively. The generated data points are green. In (a), we contrast the two SVM models looking for samples for which their predictions differ and discover that this is the case in a region near the origin lacking any training points. In (b) and (c), we inquire about data points that would be considered risky by the two models using RBF and cubic kernels, respectively. In (d) we start from a data sample shown in orange, and design $G$ so that it generates data points which are classified with the opposite label according to the RBF-SVM without straying too far from the original point.
Model-contrasting samples. Given two models, finding data points where their predictions disagree is illuminating either to interpret model-specific biases or to audit consistency between two different models. This is particularly useful when comparing models with different inductive biases (MLP vs. CNN, linear vs. nonlinear, and so on). Given $\widehat{y}$ as the predictor functions of the model that is being compared against (which can be non-parametric like in XGBoost), we solve (3) with the function
$G({\boldsymbol{\mathbf{x}}})=\ell_{G}(y_{{\boldsymbol{\mathbf{\theta^{*}}}}}({\boldsymbol{\mathbf{x}}}),1-\widehat{y}({\boldsymbol{\mathbf{x}}}))+R_{G}({\boldsymbol{\mathbf{x}}}).$ (5)
In Figure 2(a) we contrast two SVM models with different kernels and discover a region near the origin, which does not contain any samples from the dataset but would give conflicting predictions if a new sample were to come from it.
Prediction-risky samples. In order to identify the indecisive regions in a model’s decision surface, we put $G({\boldsymbol{\mathbf{x}}})=\|f({\boldsymbol{\mathbf{x}}},{\boldsymbol{\mathbf{\theta}}}^{*})-{\boldsymbol{\mathbf{\alpha}}}\|_{r}^{r}$ for $r\geq 1$, where $f$ may be the decision function of a model such as in SVM and logistic regression, or it may be the softmax probabilities of a neural network. By solving (3), we thus seek data points with probabilities close to ${\boldsymbol{\mathbf{\alpha}}}$. For binary classification, where $\alpha$ denotes the cut-off point, this corresponds to generating a “risky data point” near the decision boundary. This probing strategy can yield insights into whether the model correctly identifies important aspects of the data for decision-making and whether that aligns with the user’s conceptions. As an illustrative example in Figure 2(b), we search for the decision boundary of a RBF-SVM trained on a dataset of two concentric circles, and we discover the ring in the center as an unexpected decision boundary.
Parameter-sensitive samples. Given a set of parameters ${\boldsymbol{\mathbf{\theta}}}^{*}$ and a distribution $q^{*}({\boldsymbol{\mathbf{\theta}}})$ of parameter values, we ask the model for data samples whose classifications would flip if the model parameters were to be (perhaps slightly) perturbed. This can be achieved by solving (3) using the function
$G({\boldsymbol{\mathbf{x}}})=\int_{\Theta}\ell_{G}(y_{{\boldsymbol{\mathbf{\theta}}}}({\boldsymbol{\mathbf{x}}}),1-y_{{\boldsymbol{\mathbf{\theta}}}^{*}}({\boldsymbol{\mathbf{x}}}))q^{*}({\boldsymbol{\mathbf{\theta}}})\mathrm{d}\,{\boldsymbol{\mathbf{\theta}}}+R_{G}({\boldsymbol{\mathbf{x}}}).$ (6)
This integral would be approximated by samples from $q^{*}$. When $q^{*}\in\mathcal{Q}$ is chosen from a restricted family of distributions, like Gaussians with fixed variance, sampling from $q^{*}$ means perturbing ${\boldsymbol{\mathbf{\theta}}}^{*}$. This probing function is particularly useful when we want to examine the consistency of a model’s predictions under small shifts in its parameters. By identifying inputs whose predictions vary significantly with minor parameter changes, we can highlight sensitive regions in the input space—areas that might indicate over-dependence on specific parameter configurations. This can have overlaps but is distinct from prediction-risky samples, as we showcase in our computational study section. Parameter-sensitive samples has the flexibility to be generated far from the decision boundary, especially in non-linear models.
Fixed-label samples. Finally, we probe the model for what it thinks are good data samples that fit the bill for the prediction $y^{\prime}$, either for a single parameter ${\boldsymbol{\mathbf{\theta}}}^{*}$ or a distribution $q^{*}({\boldsymbol{\mathbf{\theta}}})$
$G({\boldsymbol{\mathbf{x}}})=\ell_{G}(y_{{\boldsymbol{\mathbf{\theta}}}^{*}}({\boldsymbol{\mathbf{x}}}),y^{\prime})+R_{G}({\boldsymbol{\mathbf{x}}})\quad\text{ and }\quad G({\boldsymbol{\mathbf{x}}})=\int_{\Theta}\ell_{G}(y_{{\boldsymbol{\mathbf{\theta}}}}({\boldsymbol{\mathbf{x}}}),y^{\prime})q^{*}({\boldsymbol{\mathbf{\theta}}})\mathrm{d}\,{\boldsymbol{\mathbf{\theta}}}+R_{G}({\boldsymbol{\mathbf{x}}}),$ (7)
respectively. Here, $R_{G}({\boldsymbol{\mathbf{x}}})$ is a localizer at an anchor point. We can take a data point $({\boldsymbol{\mathbf{x}}}_{0}.y_{0})$ to be this anchor. In case $y^{\prime}\neq y_{0}$, we are exploring changes in ${\boldsymbol{\mathbf{x}}}_{0}$ that would need to happen for the prediction to change; in other words, a counterfactual; see Figure 2(d). In case $y^{\prime}=y_{0}$ with a weak localizer, we can obtain a sample that would lead to a similar prediction, i.e., a factual.
Figure 1 demonstrates the steps when $y_{{\boldsymbol{\mathbf{\theta}}}}({\boldsymbol{\mathbf{x}}})={\boldsymbol{\mathbf{x}}}^{\intercal}{\boldsymbol{\mathbf{\theta}}}$ corresponds to linear regression, and both $\ell_{F}$ and $\ell_{G}$ are the mean squared errors. For this special case, we obtain analytical solutions for all steps of our framework. The details of this observation are given in Appendix A.

### v3

# Generating Samples to Probe Trained Models

@@ 1 Introduction
To respond to this inquiry, we study the implicit data distribution favored by trained models. In other words, our approach to understanding a model is based on creating samples in the data domain that the trained model considers favorable for a specific task. Unlike conventional ML pipelines that focus on static datasets and predictive accuracy, our approach enables dynamic interrogation of model behavior via investigating the answers of the model to specific questions. We show that these questions can be customized to each situation and they can be expressed mathematically through a loss function that evaluates the data based on a combination of data characteristics and model parameters. We consider the problem of understanding a model to be a more nuanced endeavor that requires exploration across multiple dimensions of questioning. This involves providing explanations, such as counterfactual (Wachter2017) or prototypical (Biehl2016) scenarios, shedding light not only on why a particular prediction was made, but going beyond it as well. For instance, insights into model behavior can be gained by generating parameter-sensitive data samples. When two models showing similar performances give different predictions, also known as predictive multiplicity (Marx20), our approach can be used to generate data to systematically compare their behaviors in diverse scenarios. These custom questions, and others, provide a qualitative understanding of the model. In addition, users have the flexibility to customize queries by designing specific probing functions within the data space.
@@
Related Literature. Our work complements extensive research in synthetic data generation that has been pivotal in addressing fairness, bias detection (Kusner17) and reduction (Xu18; Breugel21) as well as dataset augmentation (Wong16; Fawaz18). Generative Adversarial Networks (GANs) and Variational Autoencoders (VAEs) have been widely used to approximate data distributions (Goodfellow14; Xu18; Kingma14; Breugel24), focusing on privacy, diversity, and fidelity as primary goals.
@@
Recent studies leveraged generative models for counterfactual generation and exploring underrepresented data regions. For example, Joshi19 proposed a framework for generating task-specific synthetic data, enhancing model explainability. Similarly, Redelmeier24 introduced an approach using autoregressive generative models to create counterfactuals, facilitating bias exploration and decision boundary analysis. Recent work on global counterfactual explanations has further expanded the scope of interpretability by targeting group-level understanding. rawal2020 introduced a framework for generating global, rule-based recourse summaries for population subgroups, optimizing objectives like accuracy, coverage, and cost. These summaries are ‘if-then’ rules linking subgroup characteristics to actionable feature changes that influence a model’s prediction. Plumb20 proposed interpreting clusters in low-dimensional representations by finding sparse transformations that align one group with another.
@@
Energy-based models (EBMs) have also emerged as a promising framework, combining generative and discriminative modeling tasks. By treating classifier logits as an energy function, EBMs can model joint distributions over data and labels (LeCun06; Duvenaud20). Applications of EBMs include adversarial robustness, out-of-distribution detection, and data augmentation (Zhao17; Liu20; Arbel21; Margeloiu24). For instance, Duvenaud20 demonstrated improved out-of-distribution detection using a joint energy-based model, while Ma24 extended EBMs to tabular data for synthetic data generation.
@@
The proposed framework draws inspiration from these works while introducing a distinct perspective. Our probing function can be seen as an energy function and leads to Gibbs distribution. However, rather than learning the energy function to capture the data distribution (conditioned on label), we create a probing function using trained models. This design allows the distribution to generate samples that address the specific posed question. Related works, such as (Duvenaud20) and (Ma24), adopt a similar approach by utilizing a trained classifier to obtain an energy function and using Langevin dynamics for sampling from the Gibbs distribution. However, their main objective is to mimic the true data distribution. In fact, the former paper combines training of the energy function and classifier. In contrast, we propose a flexible framework that allows for directing diverse queries to trained models via probing functions that reflect various objectives, such as identifying prediction-risky, parameter-sensitive, or model-contrastive data samples.
Contributions. We introduce a new inductive approach that generates data samples through a flexible probing function designed to analyze and reveal the behavior of a trained model. Our method can be tailored to suit various classification and regression tasks, demonstrating its versatility in producing data that meet specific queries. This work serves as a foundational step in establishing the effectiveness and potential of our approach.
@@ 2 The Mathematical Framework
First, we present our notation. The labeled data lie in $\mathcal{X}\times\mathcal{Y}$, and the model defines a predictor function $f({\mathbf{\bm{\theta}}},\cdot):\mathcal{X}\to\mathcal{Y}^{\prime}$ for any given set of model parameters ${\mathbf{\bm{\theta}}}\in\Theta$. For a given sample ${\mathbf{\bm{x}}}\in\mathcal{X}$, the predicted label $y_{{\mathbf{\bm{\theta}}}}({\mathbf{\bm{x}}})\in\mathcal{Y}$ is obtained from the predictor function. The cost function $\ell_{F}:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$ measures how far the predicted labels are from the true labels.
The standard construction of the parameter loss function is
$F({\mathbf{\bm{\theta}}})=\int_{\mathcal{X}\times\mathcal{Y}}(\ell_{F}(y_{\bm{\theta}}(\mathbf{x}),y)+R_{F}(\bm{\theta}))\mathrm{d}\nu(\mathbf{x},y)=\frac{1}{N}\sum_{i=1}^{N}\ell_{F}(y_{{\mathbf{\bm{\theta}}}}({\mathbf{\bm{x}}}_{i}),y_{i})+R_{F}({\mathbf{\bm{\theta}}}),$ (1)
which can be seen as an integral of $\ell_{F}+R_{F}$ against the empirical distribution given by the training dataset $\{({\mathbf{\bm{x}}}_{i},y_{i})\}_{i=1}^{N}\subseteq\mathcal{X}\times\mathcal{Y}$. Here, $R_{F}({\mathbf{\bm{\theta}}})$ is a regularizer term that depends only on $\mathbf{\bm{\theta}}$}.
Figure 1 provides an overview of our framework. Just as the training process, which uses $F$ to find the right parameters, our framework probes the model with a function $G$ defined on the data space. In the variational setting, the symmetry is clear, where we get a distribution over the parameters (data) instead of a single $\mathbf{\bm{\theta}}$$^{*}$ ($\mathbf{\bm{x}}$$^{*}$). The loss function $F$ in equation 1 is an average over observed data, and, similarly, we construct the data loss $G$ by integrating out $\mathbf{\bm{\theta}}$} from a curated function (described later) that depends both on data and model parameters. We design this function of data and parameters to attain low values when the desiderata of our probing scenario are met. Different choices correspond to posing different questions to the model, and allow us to gain valuable insights into its behavior by observing the generated data, both qualitatively and through population-level statistics.
@@
Figure 1: (Left) Overview of model probing by data generation. Samples from $p^{*}({\mathbf{\bm{x}}})$ answer the question posed by $G$. The vertical arrows (a) and (c) start with functions and lead to distributions on the same space by solving equation 2 and equation 3. The diagonal arrow (b) starts with a distribution on the parameter space and obtains a loss function on the data space by integrating out $\mathbf{\bm{\theta}}$} dependence of a function on $\Theta\times\mathcal{X}$ against the distribution $q^{*}({\mathbf{\bm{\theta}}})$. (Right) The special case of the Linear Regression (LR) model with mean square error admits an analytic solution. The $G$ function is designed to find data points $\mathbf{\bm{x}}$} whose solutions under LR are close to a chosen prediction $y^{\prime}$ and averaged over $q^{*}$. The distribution $p^{*}({\mathbf{\bm{x}}})$ is calculated to be a Gaussian distribution centered at a point $\hat{{\mathbf{\bm{f}}}}$, which is shifted from the mean of given data by a certain amount depending on the desired output value $y^{\prime}$. Explicit forms of $\hat{{\mathbf{\bm{f}}}},\Sigma$ and $\widehat{{\mathbf{\bm{\theta}}}}$ and their derivation can be seen in Appendix A.
The blue arrows (a) and (c) in Figure 1 map (loss) functions on the respective spaces $\Theta$ (parameter space) and $\mathcal{X}$ (data space) to distributions over the same spaces. This corresponds to solving the Bayesian Learning Problem (BLP), which –in the case of (a)– is
@@
where $\mathcal{Q}$ is a choice of candidate distributions on $\Theta$, and $\mathcal{H}(q)=-\int_{\Theta}q\log q\,\mathrm{d}\mu$ is the entropy with respect to a base measure $\mu$. The problem can be interpreted as an implementation of the exploration-exploitation trade-off in the parameter space. The constant $\tau>0$ is called the temperature and balances these two objectives. If $\mathcal{Q}$ is the set of all density functions, then the Gibbs-Boltzmann distribution $q^{*}({\mathbf{\bm{\theta}}})\propto e^{-\frac{1}{\tau}F({\mathbf{\bm{\theta}}})}$ is the unique solution to equation 2.
Symmetrically on the data space $\mathcal{X}$, the blue arrow labeled (c) in Figure 1 represents solving
@@
The distribution $p^{*}({\mathbf{\bm{x}}})\propto e^{-\frac{1}{\tau}G({\mathbf{\bm{x}}})}$ is its global solution, balancing the expectation term’s effect of mass concentration at low $G$-values, with the entropy term’s effect of exploring the data space.
There are various methods of sampling from equation 3. In this work, we used Metropolis Adjusted Langevin Algorithm (MALA) to sample directly from the Gibbs-Boltzmann distribution $p^{*}({\mathbf{\bm{x}}})\propto e^{{-\frac{1}{\tau}G({\mathbf{\bm{x}}})}}$. This method is a kind of noisy gradient descent, with an acceptance/rejection step ensuring that the limiting distribution is $e^{-\frac{1}{\tau}G({\mathbf{\bm{x}}})}$. Details of this method are given in Appendix C. Alternatively restricting the problem to a statistical manifold $\mathcal{P}$, Variational Inference (VI) can be used to effectively reach a distribution $p^{*}\in\mathcal{P}$; see (gangulyVI; picarditer). For example, if $\mathcal{P}$ were chosen to be the Gaussians, then one would only need to keep track of the mean and the covariance in learning a $p^{*}\in\mathcal{P}$.
The red arrow (b) constructs the function $G$ as an integral of a function over both the data and parameter spaces, obtained by integrating out $\mathbf{\bm{\theta}}$} with respect to the measure $q^{*}({\mathbf{\bm{\theta}}})$, in direct analogy with the construction of $F$ in equation 1. This construction of $F$ is an integral against the empirical data distribution and, therefore, the learned parameters are compatible with the training data. Analogously, constructing $G$ ensures that the search over $\mathcal{X}$ remains compatible with parameters sampled from $q^{*}({\mathbf{\bm{\theta}}})$. In particular, if $q^{*}({\mathbf{\bm{\theta}}})$ is a is a Dirac-delta distribution, this reduces to a single parameter vector ${\mathbf{\bm{\theta}}}^{*}\in\Theta$. The specific choice of the integrand for $G$ determines which samples from $p^{*}({\mathbf{\bm{x}}})$ are the data points that answer a question posed about the trained model.
Lastly, note that we can change our search space by replacing $G$ with $G\circ\varphi$ for some $\varphi:\mathcal{Z}\to\mathcal{X}$. In high-dimensional data spaces $\mathcal{X}$, we will use this setup with $\mathcal{Z}$ as the latent space and $\varphi$ as the decoder function of a pre-trained Variational AutoEncoder (VAE). In this case, $p^{*}({\mathbf{\bm{z}}})$ becomes a distribution on $\mathcal{Z}$, and mapping its samples to $\mathcal{X}$ by $\varphi$, gives points on the data manifold. Using a pre-trained VAE for image models reduces the data space dimension, and therefore, the efficiency of the MALA sampling process, but there are also conceptual benefits. There are vast regions of the high-dimensional input space that do not correspond to plausible images and which the model did not encounter during the training process. Some of these regions may also satisfy our query $G$. Therefore, using a pre-trained VAE decoder reflects a plausible image requirement, if imposed.
@@ 3 Probing Trained Models
We start with a general structure of the loss function curated for questioning trained models:
$G({\mathbf{\bm{x}}})=\int_{\Theta}\ell_{G}(y_{{\mathbf{\bm{\theta}}}}({\mathbf{\bm{x}}}),\widehat{y}({\mathbf{\bm{x}}}))q^{*}({\mathbf{\bm{\theta}}})\mathrm{d}\,{\mathbf{\bm{\theta}}}+R_{G}({\mathbf{\bm{x}}}),$ (4)